OpenStudy (anonymous):

what points lies on the circle whose center is at the origin and whose radius is 13?

OpenStudy (anonymous):

x^2+y^2=169

OpenStudy (anonymous):

by formula (x-a)^2+(y-b)^2=r^2

OpenStudy (anonymous):

(a,b) is center r= radius

OpenStudy (anonymous):

(-5, 12) (sqrt 13,sqrt13 ) (6, -7)